Consider a call option with strike price $100 on a stock. There are two periods until the option expires. The stock can either move up by u = 1.25 or down by a factor d = 0.8 in each period. Assume the interest rate for each period is given by r̂ = e^r with r = 5% and that the current stock price is $95. Find the value of the call option.

(Please write quick explanation for each step of the process, thank you so much!)
Answer #1

Risk neutral probability of up state = p = (er - d) / (u - d) = (e0.05 - 0.8) / (1.25 - 0.8) = 0.55838

Node S(T) Payoff Probability Probability Probability x payoff
max(S - K, 0) = max (S - 100,0) Formula Value
Suu                  148.44                                             48.44 p2 0.311788464                           15.10
Sud                    95.00                                                   -   2p(1 - p) 0.493183501                                 -  
Sdd                    60.80                                                   -   (1-p)2 0.195028035                                 -  
Total 1                           15.10

Hence, value of the call option today = 15.10 x e-r x t = 15.10 x e-0.05 x 2 = $  13.67